The following table shows the number of males in a country in 2000, broken down by age. Numbers are in millions.

Age 0–18 18–25 25–35 35–45 45–55 55–65 65–75 75
and
over
Number 36.9 15.5 21.4 22.4 18.1 12.3 8.1 6.1


Construct the associated probability distribution (with probabilities rounded to four decimal places).

Age 0–18 18–25 25–35 35–45 45–55 55–65 65–75 75
and
over
Probability


Use the distribution to compute the following. (Round your answers to four decimal places.)

(a)  P(25 ≤ X ≤ 65)

(b)  P(X ≤ 18)

(c)  P(X ≥ 15)
Answer #1
Age Prob.
0- 18 0.2621
18-25 0.1101
25-35 0.1520
35-45 0.1591
45-55 0.1286
55-65 0.0874
65-75 0.0575
75 and 0ver 0.0433

P(25< = x < 65)
= P(x< =65) - P(x< =25)
= 0.1520 + 0.1591+ 0.1286+0.0874
= 0.5271


b)

P(x< =18)
= 0.2621

c)

P(x>= 15) = 1